Bramble Industries carries no inventories. Its product is manufactured only when a customer’s order is received. It is then shipped immediately after it is made. For its fiscal year ended October 31, 2017, Bramble’s break-even point was $1.50 million. On sales of $1.40 million, its income statement showed a gross profit of $280,000, direct materials cost of $400,000, and direct labor costs of $508,000. The contribution margin was $280,000, and variable manufacturing overhead was $49,000.

(a) Calculate the following: (Round intermediate calculations to 2 decimal places e.g. 2.25 and final answers to 0 decimal places, e.g. 1,225.)

1. Variable selling and administrative expenses. $

2. Fixed manufacturing overhead. $

3. Fixed selling and administrative expenses. $


(b) Ignoring your answer to part (a), assume that fixed manufacturing overhead was $102,000 and the fixed selling and administrative expenses were $76,000. The marketing vice president feels that if the company increased its advertising, sales could be increased by 21%. What is the maximum increased advertising cost the company can incur and still report the same income as before the advertising expenditure?

Maximum increased advertising expenditure $
Answer #1

PART A (1)

Variable Selling and Administrative Expenses = Sales - Direct Material Costs - Direct Labor Costs - Variable Manufacturing Costs - Contribution Margin

Variable Selling and Administrative Expenses = 1,400,000 - 400,000-508,000-49,000-280,000 = $ 163,000

PART A (2)

Fixed Manufacturing Overhead = Sales - Direct Material Costs - Direct Labor Costs - Variable Manufacturing Costs - Gross Profit

= 1,400,000 - 400,000-508,000-49,000-280000 = $ 163,000

PART A (3)

Contribution Ratio = Contribution*100/Sales = 280,000*100/1,400,000 = 20%
Total Fixed Cost = Break Even*Contribution Ratio = 1,400,000*20% = $280,000
Fixed selling and administrative expenses = Total Fixed Cost - Fixed Manufacturing overheads = 280,000 - 163,000 = $117,000

PART B

Maximum Increased Revenue Expenditure = (Current Revenues and Expenses)*(1+Revenue Increase Percentage) - Manufacturing Overhead = (102,000 + 76,000)*(1 + 21%) - 102,000 = $113380

Hope this helps, if not please let know in comments. We expert here put lots of effort to give correct and quality answer to students. Please mark the answer as helpful for the efforts put, it will mean alot. Thanks